An unnamed company is designing a product layout for a new product. It plans to use this production line to meet projected demand of 320 units per 7.00 hour day, and the tasks necessary to produce this product are as follows...

Task Time (sec) Predecessor
u 30 none
v 33 u
w 4 v
x 17 u, v
y 51 x
z 23 v, y

If the company wants the output rate to be equal to demand, what is the desired cycle time in seconds?
Answer #1

Total Task Time = 30+33+4+17+51+23 = 158 seconds

Total Available Time = 7*60*60 = 25200 seconds

Output rate = Demand = 320 units per day

Desired Cycle Time = Total Available Time/Demand

Desired Cycle Time = 25200/320 = 78.75 seconds